OpenStudy (anonymous):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0)

OpenStudy (anonymous):

@***[ISURU]*** @littlelotte @ganeshie8 @jim_thompson5910 @Philoctetes Help me plz

OpenStudy (anonymous):

y - 3x=7 21 + 9x = 3y first y= 3x + 7 then substitute 21 + 9x = 3 (3x +7) 21 = -9x + 9x + 21 0 = 0 i think there is no solution